help me to complete the table based on the information given about the element Boron.

Number Element    Atomic Number    Mass Number Atomic mass

1. Boron with 5

neutrons

2. Boron with 6

neutrons

3. Boron with 7

neutrons

4. Boron with 8

neutrons
Answer #1

Atomic number of boron is 5. All the elements are Boron's isotopes:

1)Atomic number of boron is 5. Atomic mass = 5+5 = 10

2) Atomic number of Boron is 5. Atomic mass = 5+6 = 11

3) Atomic number 5 and atomic mass = 5+7 = 12

4) atomic number 5 and atomic mass = 5+8 = 13